OpenStudy (anonymous):

find the maximum y value f(x)=-x^2+6x+9 please help me solve

OpenStudy (anonymous):

3

OpenStudy (jziggy):

Using derivatives: f'(x) = -2x + 6 Solve for when it equals zero -2x + 6 = 0 -2x = - 6 x = 3